Consider the probability that at most 81 out of 476 computers will crash in a day. Choose the best description of the area under the normal curve that would be used to approximate binomial probability.
Answer #1

We need to find the probability that at most 81 out of 476 computers will crash in a day.

P(X \le 81)

Using continuity correction, we know that

P(X \le A) \rightarrow P(X<A+0.5 )\\ Or\ P(X \le 81) \simeq P(X<81.5)

Less than sign indicates that the area under curve is to the left side of 81.5

Therefore, area under the normal curve that would be used to approximate binomial probability is area to the left of 81.5